There is a 30% chance of rain today. If it does not rain today, there is a 20% chance of rain tomorrow. If it rains today, there is a 50% chance of rain tomorrow. What is the probability that it rains tomorrow?
1
Expert's answer
2020-03-02T15:08:46-0500

Chance that it doesn't rain today and it will rain tomorrow: (1-0.3)*0.2=0.14

Chance that it's raining today and it's going to rain tomorrow: 0.3*0.5=0.15

The chance that it will rain tomorrow: 0.14+0.15=0.29 or 29%